OpenStudy (anonymous):

An electric field is measured at 3.00 × 10^5 N/C at a distance from a source point charge of 3.00 × 10^-6 C. How far is the point from the source charge? A. 0.09 m B. 0.30 m C. 0.36 m D. 0.60 m

OpenStudy (michele_laino):

the requested distance is given by the subsequent formula: \[\large d = \sqrt {\frac{{KQ}}{E}} = \sqrt {\frac{{9 \times {{10}^9} \times 3 \times {{10}^{ - 6}}}}{{3 \times {{10}^5}}}} \]

OpenStudy (anonymous):

ok! so we get 0.3? so choice B is our solution?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay!! thank you!!:D

OpenStudy (michele_laino):

:)